You have an investment account that started with

​$4 comma 0004,000

1010

years ago and which now has grown to

​$9 comma 0009,000.

a. What annual rate of return have you earned​ (you have made no additional contributions to the​ account)?

b. If the investment account earns

15 %15%

per year from now​ on, what will the​ account's value be

1010

years from​ now?

a. What annual rate of return have you earned​ (you have made no additional contributions to the​ account)?

Your annual rate of return is

nothing​%.

​(Round to two decimal​ places.)b. If the investment account earns

15 %15%

per year from now​ on, what will the​ account's value be

1010

years from​ now?The​ account's value in ten years will be

​$
Answer #1

1.
Annual rate of return=(9000/4000)^(1/10)-1=8.44717711976986%

2.
Account's value=9000*(1+15%)^10=36410.0196213712